OpenStudy (anonymous):

Whilst at University, Oswald's aunt supported him financially. On 1 July 2012 Oswald agreed to pay her $10000 on 1 July 2017 and a further $10000 on 1 July 2022. At current rate of interest, the present value of these payments equal to $18762.47. What is the duration of Oswald's liabilities to his aunt?

OpenStudy (tkhunny):

Third time I've seen this one. Is no one teaching "Basic Principles" out there? $10000v^5 + $10000v^10 = $18762.4 Can you solve for v?

OpenStudy (anonymous):

how do i find duration in years after found i? thx

OpenStudy (tkhunny):

You'll need a definition of "Duration". Do you have one? Please share.

OpenStudy (anonymous):

Definition of Duration: For some holding period the rise in price is exactly offset by the decrease in return from the reinvested coupons , This special holding period is known as the Duration of the bond So how can i work out thx

OpenStudy (tkhunny):

Well, okay, there is some ambiguity, here. It is the nature of the word "Duration". You have supplied the defintion of "Modified Duration". For this definition of "Duration", we'll have to decide on what "Unit Change" to use for the interest rate. Gererally, 1% is a very large change. Why not use 1 Basis Point, or 1% of 1%. Once you solve this for i, $10000v^5 + $10000v^10 = $18762.4, calculate the new price using i + 0.0001 and see what happens. You may also want to calculate with i - 0.0001 to check for symmetry. Sound like a plan?

OpenStudy (anonymous):

quite make sense im working on it

OpenStudy (anonymous):

The i is 0.00858355 that I cal SO using the hint u told me I figure out the D =8.6667 years I cant desire to ask u the final answer but its make sense? thx so much!

OpenStudy (tkhunny):

Well, there's that ambiguity again. The "Duration" you have defined is NOT measured in years. If you want to calculate a "McCauley Duration", you can report it in years.

OpenStudy (anonymous):

I actually calculator by finding Pi1 and Pi0 Using the MD = D/(1+i) I think I got it thanks again!

OpenStudy (tkhunny):

Fair enough. Any time anyone just says "Duration", don't let them get away with it. Make them tell you what they mean.

OpenStudy (anonymous):

I pretty sure for this Question. I was wondering if u could help me with the one Floting rate bond Question? Its quite urgeny pls help

OpenStudy (anonymous):

Once again thx!